A system does 2.90 ✕ 108 J of work while 9.40 ✕ 108 J of heat transfer occurs to the environment. What is the change in internal energy of the system assuming no other changes (such as in temperature or by the addition of fuel)? (Include the sign of the value in your answer.)

J
Answer #1

according to thermodynamic first law

Q = U+W

Q = HEAT

U = INTERNAL ENERGY

W = WORK

U = Q-W = (9.4-2.9)*10^8 = 6.5*10^8 J